OpenStudy (anonymous):

-1/4y=16 a.-64 b.64 c.4 d.-4

jigglypuff314 (jigglypuff314):

to solve for y you would want to get y by itself to do that, you would have to multiply both sides by the reciprocal of -1/4 therefore you would multiply both sides by -4 so y = 16*(-4) =

OpenStudy (anonymous):

so -4 is the answer

jigglypuff314 (jigglypuff314):

umm, not quite... 16 * 4 = 64 so 16 * (-4) = -__

OpenStudy (anonymous):

-64

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thanks

jigglypuff314 (jigglypuff314):

you're welcome :)